# Observation of Kolmogorov turbulence due to multiscale vortices in dusty plasma experiments

### v1

@@ II.2 Image processing and PIV
Figure 5: Time-elapsed images (20 consecutive frames for each case) of dust rotation in campaigns I and II. Their locations and discharge parameters are as per Table  1.
@@ III Results and Discussions
We report the observation of self-excited dust rotation dynamics in the 3D dust cloud within a weak plasma. We analyze its properties to understand its turbulence characteristics. The following is a detailed discussion of the results.
Figure 6: Time-elapsed images (20 consecutive frames for each case) of dust rotation in campaigns III-VI. Their locations and discharge parameters are as per Table  1.
@@ III.1 Dust vortex formation
Figure 7: Vorticity profile for dust vortex flow in case of campaigns I and II. The dust cloud formed above the glass surface at low discharge voltages.
Figure 8: Vorticity profile for dust rotation in regions III-VI of the dust cloud formed above the glass surface at high discharge voltages.
The dust cloud in the diffused plasma region showed a variety of collective dynamics as we varied the discharge voltage between 230 and 600 volts. As listed in table 1, the self-sustained dust rotation has been observed for values at both edges of the voltage range. Figures 5 and 6 show time-elapsed images for different campaigns where dust rotation is visible from particle trajectories. The exact location of each dust cloud with reference to the center of the cathode (in Z-direction) and the surface of the glass (in X-direction) is mentioned as axis labels. All other dust rotations are observed typically at the same location and similar discharge parameters except for campaigns I and II. We have provided dust rotation movies for each campaign as supplementary material mov .
@@
A comprehensive analysis to establish the cause of dust rotation is beyond the scope of this paper. However, we ruled out some of the possibilities. A temperature gradient in neutral gas can give rise to thermophoretic force which can affect dust dynamics (Shukla and Mamun, 2015). Typically in glow discharge, the temperature of the cathode can increase reasonably due to ion bombardment after a few hours of operation (Kaur et al., 2015b). In our experiment, we observed the dust vortex soon after reaching the operating discharge voltage range before the cathode was reasonably heated. This suggests that thermophoretic force may not influence the vortex dynamics significantly. We also ruled out the possibility of a neutral flow in the vacuum chamber from affecting the dust dynamics. We repeated the vortex experiment by simultaneously closing the pump and the Ar gas inlet to minimize any kind of neutral flow in the vacuum chamber. In this condition, too, we observed the dust vortex, thus ruling out any role of neutral flow.
@@
Figures 7 and 8 provide flow velocity vectors and vorticity plots for all campaigns. The vorticity plots were obtained after taking the mean of velocity field profiles from 100-125 frames, which profoundly helps visualize the large-scale rotations. However, it washes out the information on small-scale vortex formations, which is important when evaluating the energy spectrum in the next section. The vorticity profiles of all campaigns reflect that the dust rotation is clockwise for campaigns III-VI as shown in Fig. 8, and the same is counterclockwise for campaigns I and II (see Fig. 7). This is observable from the velocity flow vectors and the colors of the vorticity profiles in different plots.
Figure 9: Time-averaged power spectrum of velocity fields obtained from PIV for campaigns I to VI with error bars corrected by a factor 0.5 to avoid log of negative values in log scale. The wavevectors are normalized with $\mathrm{k_{max}}$ to get spectra in the same scales. All spectra follow the standard Kolmogorov scaling of $-5/3$.
III.2 The spatial power-spectrum
@@
An important feature of turbulent flows is that they exhibit multi-scale energy cascades usually demonstrated through spatial energy spectrum (Kolmogorov, 1991a; Frisch and Kolmogorov, ). Figure 9 presents subplots displaying the energy spectra for each campaign, which demonstrates the distribution of kinetic energy across different scales or wavenumbers. The spectra are calculated as follows. For each case, the vortex motion is quite stable, and dust continues to rotate until we change the plasma creation conditions. Thus, we take the PIV for each video and get multiple velocity-flow profiles for each campaign for a one-second video. Once we have a coarse-grained flow velocity ${\bf u(x},t)$ for a given snapshot, we perform a Fourier transform on ${\bf u(x},t)$ to obtain ${\bf u(k},t)$, from which we calculate the modal energy $E({\bf k},t)=|{\bf u(k},t)|^{2}/2$. Afterward, we determine the one-dimensional shell spectrum $E(k,t)$ by summing the modal energies over a specific range $k-1\leq k^{\prime}\leq k$ as explained by Frisch and Kolmogorov .
Figure 10: The second-order structure functions $S_{2}(l)$ are shown for six experiments (I-VI). Dashed lines indicate the theoretical $l^{2/3}$ scaling predicted by Kolmogorov’s theory in the inertial range for reference. Each plot presents results from different experimental setups, highlighting variations in inertial range scaling and the extent of Kolmogorov scaling.
@@
The power spectra were independently estimated for each flow profile of the recorded datasets, and the average energy spectra over multiple time frames were then plotted with error bars, as shown in Figure 9 for all the campaigns. The wavevectors are normalized with $\mathrm{k_{max}}$, where $\mathrm{k_{max}}$ is the maximum wavenumber associated with the system. The values of $\mathrm{k_{max}}$ for each campaign are $6.334$, $5.69$, $5.69$, $5.69$, $4.91$, and $4.91$ $\mathrm{mm^{-1}}$, respectively. Note that the units of $\mathrm{k_{max}}$ is $\mathrm{mm^{-1}}$, and the units of the energy spectra are $\mathrm{(mm/s)^{2}}$. The normalization is performed to present the spectra on the same scale without losing generality.
@@
We observe that the energy fed at the large scales cascades to the smaller scales following the standard -5/3 scaling for all campaigns, which is consistent with Kolmogorov’s theory (Kolmogorov, 1991b). The corresponding scaling in the real space is observed by analyzing the second order structure function discussed in section III.3. However, it should be noted that the wavevector ranges vary slightly for different campaigns. Also, the total energy range in the spectrum varies up to an order of magnitude as we look for the spectrum for I-VI. It is based on the vortex domain size with finite flow velocities. For example, in Fig. 7, the vortex size of I is smaller than that of II; hence, its energy spectrum has a larger magnitude range.
@@ III.3 Second-order structure factor
We analyze the structure-function for examining the statistical characteristics of turbulent flows (Frisch and Kolmogorov, ). Structure functions represent the moments of velocity increments $\delta u^{p}_{r}=\langle[u(\vec{r}+l)-u(\vec{r})]^{p}\rangle$, where $r$ denotes a reference point, $p$ is the order of the structure-function, and $l$ represents the separation distance between two points within a velocity field (Chatterjee, 2023). In accordance with Kolmogorov’s theory (Kolmogorov 1941b) for fully developed homogeneous turbulence, it is anticipated that the scaling relationships in the inertial range of turbulence adhere to $\delta u^{p}_{r}\propto r^{\xi(p)}$, where $\xi(p)=p/3$ approximately (Benzi et al., 1993).
Figure 11: Power spectrum in frequency domain obtained by taking FFT of time correlation of velocity field. The spectrum typically follows a $-5/3$ scaling similar to that observed in the spatial domain in all experimental campaigns. The dashed line with slope $-5/3$ in each graph is for reference purposes only.
In this study, we calculate the second-order structure function for all data sets by determining the difference in velocity between a reference point $r$ and varying separation distances $l$, as shown in Figure 10. It is important to note that the structure functions were computed based on a time average of 10 velocity frames without the loss of generality. We find the exact scaling relation $S_{2}(l)\propto l^{2/3}$ between the second-order structure function and the incremental length as predicted by Kolmogorov (Kraichnan, 1974). This strong agreement with the anticipated power-law scaling supports our results. We also observe that the $S_{2}(l)$ values for campaigns III and IV are two orders of magnitude smaller due to the relatively lower velocity order in these vortex regimes, which is also reflected in the energy spectra, Fig 9.
III.4 Power spectrum of time-series
We also analyzed the power spectrum in the frequency domain, which gives energy distribution across various frequencies (wavelengths). This is done by taking the Fourier transform of velocity autocorrelation in accordance with the Wiener-Khinchan theorem  (Frisch and Kolmogorov, )
@@ III.4 Power spectrum of time-series
The velocity time series is computed by taking the mean of velocities of a few (around 5) velocity elements in each frame at a particular location, which typically falls on the dust rotation regime. Subsequently, we followed the above methodology to compute the power spectrum. We observed a Kolmogorov $-5/3$ power law over an initial range of frequencies as demonstrated in figure 11. The spectra were computed for all six campaigns, and the scaling is consistent across all cases.
@@
Figure 12: The Probability Distribution Function (PDF) of velocities and its gradients for campaigns (I) and (II). These PDFs are fitted with the corresponding Gaussian distribution function, represented by the black dashed lines. Note that the PDFs of $v_{z}$ and $\delta v_{z}$ are adjusted by a multiplier of 0.001 for better visualization.
@@ III.5 Probability distribution functions
Since the turbulent velocity fields exhibit significant fluctuations, it’s essential to comprehend their statistical properties. Consequently, researchers have focused on studying velocity fields’ Probability Distribution Functions (PDF) and their derivatives. Velocity PDFs provide insights into the distribution of flow velocities at different points in space (Pandit et al., 2009; Andrés and Banerjee, 2019). In fully developed turbulent flows, the PDF of velocity fields is anticipated to adhere to a Gaussian distribution, whereas their derivatives often deviate from Gaussian behavior, particularly at the tails (Gulitski et al., 2007).
Figure 12 shows the PDF of velocity gradients for all six experimental datasets. The PDF of each velocity field component $v_{z}$ and $v_{x}$ shows a Gaussian distribution with an approximate zero mean value. The PDF of velocity gradients deviates strongly from the Gaussian distribution at the tails, which is attributed to the intermittency effects.
@@
In this work, we clearly demonstrated the presence of Kolmogorov turbulence in a three-dimensional dusty plasma. The dust cloud formed naturally as a result of the forces arising in plasma and device configuration. The energy spectrum scaling of -5/3 and second-order structure factor with a slope of 2/3 are reproducible for different campaigns regardless of the origin and location of the vortices. We could also demonstrate the characteristics of the stationary turbulence due to the presence of -5/3 scaling in both spatial and time-series spectra. The self-consistent nature of the vortex motion of charged dust clouds also points to constant energy pumping up at the largest scales and dissipated as dust’s thermal motion. Overall, the dusty plasma medium acts as a driven-dissipative system.
@@
An important aspect of the present work is the demonstration of fully developed turbulence. Traditional understanding suggests the possibility of fully developed turbulence at a high Reynolds number $R_{e}$ (Kolmogorov, 1991c). Turbulence in dusty plasmas occurs at low Reynolds numbers (Tsai et al., 2014; Zhdanov et al., 2015) similar to that in viscoelastic fluids. Our calculations suggest typical values of $R_{e}$ between $8$ and $50$. The Section V appendix includes the calculation of the Reynolds number. Examining turbulence through multi-scale vortex dynamics in a dusty plasma at such low Reynolds numbers is an important phenomenon. Although turbulence in dusty plasma experiments has been attempted in these low Reynolds number regimes, the proper energy transfer mechanism in dusty plasma experiments is still missing. Though dusty plasmas are known for their viscoelastic properties, we require a comprehensive analysis of flow before clarifying our results as elastic turbulence. This will also require establishing the $-5/3$ scaling for non-Newtonian fluid turbulence.
Figure 13: The sixth-order structure functions $S_{6}(l)$ are shown for six experiments (I-VI). Dashed lines indicate the scaling which deviates from expected $l^{2}$ due to intermittency.
V Appendix
V.1 Calculation for Reynolds number, $R_{e}$
We calculated the Reynolds number, $R_{e}$, for our experiments using the following parameters: $\Gamma=1500$, $\kappa=1.5$, $m=6\times 10^{-13}$ kg, $n=2\times 10^{10}$ m-3, $\rho=1.2\times 10^{-2}$, $a=250\mu$m, and $\omega_{pd}=60$ Hz. From the literature, we took the normalized dynamic shear viscosity for 3D dusty plasma to be 0.2  (Donkó and Hartmann, 2008; Saigo and Hamaguchi, 2002; Sanbonmatsu and Murillo, 2001).
The normalization factor is $\eta^{*}=mn\omega_{pd}a^{2}=4.5\times 10^{-8}$ Pa.s. Thus, the dynamic shear viscosity is $\eta=0.2\times\eta^{*}=9\times 10^{-9}$ Pa.s. Using a typical length scale $l=2\times 10^{-2}$ m and a typical velocity scale $v=1\times 10^{-3}$ m/s, we calculated the Reynolds number as $R_{e}=\rho lv/\eta$ = 26. Depending on the range of parameter values, the Reynolds number in these experiments varies between 8 and 50.
V.2 Measurement of intermittency, $S_{6}(l)$
The Kolmogorov $p/3$ scale in structure functions deviates at higher orders due to intermittency (Anselmet et al., 1984). Specifically, the sixth-order structure function captures the intermittency in turbulent flows. The scaling exponent for the sixth-order structure functions deviates from Kolmogorov’s theory as $(p/3-\mu)$, where $\mu$ is an intermittency parameter (Vindel et al., 2008; Anselmet et al., 1984).
In our study, we plotted the sixth-order structure functions as shown in Figure 13. For campaigns I, II, and V, the intermittency parameters are found to be $\mu=0.6$, $0.7$, and $1$, respectively. However, for campaigns III, IV, and VI, the parameters are $-0.5$, $-1$, and $-0.5$, for which we currently do not have a proper explanation and require further investigation.
VI Acknowledgements

### v2

@@ II.2 Image processing and PIV
Figure 5: Time-elapsed images (20 consecutive frames for each case) of dust flow in campaigns I and II. Their locations and discharge parameters are as per Table  1. The inset for I highlights the trajectories of a few selected particles within the cloud.
@@ III Results and Discussions
We report the observation of self-excited dust rotation dynamics in the 3D dust cloud within a diffused plasma. We analyze its properties to understand its turbulence characteristics. The following is a detailed discussion of the results.
Figure 6: Time-elapsed images (20 consecutive frames for each case) of dust flow in campaigns III-VI. Their locations and discharge parameters are as per Table  1.
@@ III.1 Dust vortex formation
Figure 7: Vorticity profile for dust flow in case of campaigns I and II. The dust cloud formed above the glass surface at low discharge voltages.
Figure 8: Vorticity profile for dust flow in regions III-VI of the dust cloud formed above the glass surface at high discharge voltages.
Figure 9: The time-averaged energy spectra for Campaigns I–VI have been fitted with slopes that approximately match Kolmogorov’s scaling for 3D turbulence. To present the spectra on a consistent scale, wavevectors are normalized by $\mathrm{k_{max}}$.
The dust cloud in the diffused plasma region showed a variety of collective dynamics as we varied the discharge voltage between 230 and 600 volts. As listed in Table 1, the self-sustained dust rotation has been observed for values at both edges of the voltage range. Figures 5 and 6 show time-elapsed images for different campaigns where dust rotation is visible from particle trajectories. The inset in Fig. 5(I) shows the tracking of selected particles in the box region. The particle tracking was performed using the ImageJ software’s tracking algorithm. The exact location of each dust cloud with reference to the center of the cathode (in the Z-direction) and the surface of the glass (in the X-direction) is mentioned as axis labels. All other dust rotations are observed typically at the same location and similar discharge parameters except for campaigns I and II. For each campaign, the vortex motion is quite stable, and dust continues to rotate until we change the plasma creation conditions. We have provided dust rotation movies for each campaign as supplementary material mov .
@@
A comprehensive analysis to establish the cause of dust rotation is beyond the scope of this paper. However, we ruled out some of the possibilities. A temperature gradient in neutral gas can give rise to thermophoretic force, which can affect dust dynamics (Shukla and Mamun, 2015). Typically, in glow discharge, the temperature of the cathode can increase reasonably due to ion bombardment after a few hours of operation (Kaur et al., 2015b). In our experiment, we observed the dust vortex soon after reaching the operating discharge voltage range before the cathode was reasonably heated. This suggests that thermophoretic force may not influence the vortex dynamics significantly. We also ruled out the possibility of a neutral flow in the vacuum chamber affecting the dust dynamics. We repeated the vortex experiment by simultaneously closing the pump and the Ar gas inlet to minimize any kind of neutral flow in the vacuum chamber. In this condition, too, we observed the dust vortex, thus ruling out any role of neutral flow.
@@
Figures 7 and 8 provide flow velocity vectors and vorticity plots for all campaigns. The vorticity plots were obtained after taking the mean of velocity field profiles from 100 frames, which profoundly helped to visualize the large-scale rotations. The averaging cleans up the fluctuations and reflects the large-scale vortex structures. However, these fluctuations are important when evaluating the energy spectrum in the next section. The vorticity profiles of all campaigns reflect that the dust rotation is clockwise for campaigns III-VI as shown in Fig. 8, and the same is counterclockwise for campaigns I and II (see Fig. 7). This is observable from the velocity flow vectors and the colors of the vorticity profiles in different plots.
III.2 The spatial energy spectra
@@
An important feature of turbulent flows is that they exhibit multi-scale energy cascades, which are typically demonstrated through the spatial energy spectrum (Kolmogorov, 1991a; Frisch and Kolmogorov, ). Figure 9 presents subplots displaying the energy spectra for each campaign, which demonstrate the distribution of kinetic energy across different scales or wavenumbers. The spectra are calculated in the following way: For each recorded dataset, we take the PIV and get multiple velocity-flow profiles for each campaign. Once we have a coarse-grained flow velocity ${\bf u(x},t)$ for a given snapshot, we perform a Fourier transform on ${\bf u(x},t)$ to obtain ${\bf u(k},t)$, from which we calculate the modal energy $E({\bf k},t)=|{\bf u(k},t)|^{2}/2$. After that, we find the one-dimensional shell spectrum $E(k,t)$ by adding up the modal energies within a certain range $k-1\leq k^{\prime}\leq k$, as described by Frisch and Kolmogorov .
Figure 10: Log-log plots of time-averaged second-order structure function $S_{2}(l)$ vs. incremental length $l$ for campaigns (I-VI). Each plot displays the results of different experimental setups, demonstrating variations in inertial range scaling and Kolmogorov scaling degree. The least squares fitted lines (black dashed) yield a slope with standard deviation.
@@
We estimated the energy spectra for each flow profile of the recorded datasets separately. Then, we plotted the average energy spectra over several time frames with error bars, as seen in Fig 9 for all the campaigns. The wavevectors are normalized with $\mathrm{k_{max}}$, where $\mathrm{k_{max}}$ is the maximum wavenumber associated with the system. The values of $\mathrm{k_{max}}$ for each campaign are $6.334$, $5.69$, $5.69$, $5.69$, $4.91$, and $4.91$ $\mathrm{mm^{-1}}$, respectively. Note that the units of $\mathrm{k_{max}}$ are $\mathrm{mm^{-1}}$, and the units of the energy spectra are $\mathrm{(mm/s)^{2}}$. We perform normalization to present the spectra on the same scale without losing generality.
@@
We observe that the energy fed at the large scales cascades to the smaller scales following the standard -5/3 scaling for all campaigns, which is consistent with Kolmogorov’s theory (Kolmogorov, 1991b). The corresponding scaling in the real space is observed by analyzing the second-order structure function discussed in section III.3. However, note that the wavevector ranges vary slightly for different campaigns. Also, the total energy range in the spectra varies up to an order of magnitude as we look for the spectra for I-VI. It is based on the vortex domain size and finite flow velocities. For example, in Fig. 7, the vortex size of I is smaller than that of II; hence, its energy spectrum has a larger magnitude range.
The flow dynamics is three-dimensional, but the present analysis is carried out only in the x-z plane due to the limitation of recording only one 2D plane using a laser sheet and a camera. However, we have individual movies of the y-z and x-z planes (asynchronous) for campaigns V and VI. Here, we provide in Fig. 14 (see the Appendix), the energy spectrum and the vorticity plots for campaign V in the y-z plane. These plots demonstrate the 3-D characteristics of turbulent vortex flow in the cloud at the specific position of $x=5$ mm. The least-square fitted scaling (the slope) obtained is $-1.595\pm 0.099$, consistent with our claims of Kolmogorov scaling within experimental uncertainties.
@@ III.3 Second-order structure factor
We analyze the structure-function for examining the statistical characteristics of turbulent flows (Frisch and Kolmogorov, ). The structure functions display the moments of velocity increase. $\delta u^{p}_{r}=\langle[u(\vec{r}+l)-u(\vec{r})]^{p}\rangle$, where $r$ is a reference point, $p$ is the order of the structure function, and $l$ is the distance between two points in a velocity field Chatterjee (2023). Based on Kolmogorov’s theory (Kolmogorov 1941b), for fully developed homogeneous turbulence, the scaling relationships in the inertial range of turbulence should follow the form $\delta u^{p}_{r}\propto r^{\xi(p)}$, where $\xi(p)=p/3$ roughly;(Benzi et al., 1993).
Figure 11: Log-log plots of the frequency spectra are computed by taking the FFT to the time correlation of the velocity field. The spectra show a slight deviation from the Kolmogorov $-5/3$ prediction.
In this study, we calculate the second-order structure function for all data sets by determining the difference in velocity between a reference point $r$ and varying separation distances $l$, as shown in Fig. 10. To maintain generality, we computed the structural functions using a time average of 90 velocity frames. The fitted slopes yield the scaling exponent close to $0.66$ with standard deviation, which is almost consistent with Kolmogorov theory (Kraichnan, 1974). We also see that the $S_{2}(l)$ values for campaigns III and IV are two orders of magnitude lower. This is because the velocity order is lower in these vortex regimes, which is also shown in the energy spectra, Fig 9.
III.4 The energy spectra of time-series
We also analyzed the energy spectra in the frequency domain, which gives energy distribution across various frequencies (wavelengths). This is done by taking the Fourier transform of velocity autocorrelation in accordance with the Wiener-Khinchan theorem  (Frisch and Kolmogorov, ).
@@ III.4 The energy spectra of time-series
We collect multiple velocity time series from different spatial locations, which usually fall on the dust rotation region. Subsequently, we followed the above methodology to compute the energy spectra for each time series. The mean frequency spectrum and the standard deviation have been provided for each campaign. We observed a Kolmogorov $-5/3$ power law over an inertial range of frequencies, as demonstrated in figure 11. We computed the spectra for all six campaigns and found consistent scaling across all cases.
@@
Figure 12: The PDFs of velocity gradients $\delta v_{x}$ and $\delta v_{z}$ for campaigns I-VI. The black dashed lines represent the corresponding Gaussian distribution fits. For improved visualization, we have artificially shifted the PDFs of $\delta v_{x}$ by a factor of 0.01.
@@ III.5 Probability Distribution Functions
Since the turbulent velocity fields exhibit significant fluctuations, it’s essential to comprehend their statistical properties. Consequently, researchers have focused on studying velocity fields’ PDFs and their derivatives. Velocity PDFs provide insights into the distribution of flow velocities at different points in space (Pandit et al., 2009; Andrés and Banerjee, 2019). The PDF of velocity fields should follow a Gaussian distribution in fully developed turbulent flows, but their derivatives often deviate from the Gaussian distribution, especially at the tails (Gulitski et al., 2007).
Figure 12 presents the PDFs of velocity gradients, $\delta v_{z}$ and $\delta v_{x}$, which deviate from the Gaussian distribution (shown by the black dashed lines) at the tails, indicating intermittency. Note that the PDFs shown are averaged over 90 datasets for each campaign. The fitted plots are obtained by providing $\delta v_{x}$ and $\delta v_{z}$ dataset to the standard MATLAB function “fitdist” for “normal” distribution. The distribution returns a mean value and the associated standard deviation in each case. Table S2 detailing mean velocity, standard deviation, skewness, and kurtosis is included as supplementary material.
@@
In this work, we demonstrated the presence of Kolmogorov turbulence in a three-dimensional dusty plasma. The dust cloud formed naturally as a result of plasma and device configuration forces. The energy spectrum scaling of -5/3 and the second-order structure factor with a slope of 2/3 are reproducible for different campaigns, regardless of the vortices’ origin and location. We could also demonstrate the characteristics of stationary turbulence due to -5/3 scaling in both spatial and time-series spectra. The self-consistent nature of the vortex motion of charged dust clouds also points to constant energy pumping up at the largest scales and dissipating as a thermal motion of dust particles. Overall, the dusty plasma medium acts as a driven dissipative system.
@@
An important aspect of the present work is the demonstration of fully developed turbulence. Traditional understanding suggests the possibility of fully developed turbulence at a high Reynolds number $R_{e}$ (Kolmogorov, 1991c). In dusty plasmas, earlier studies reported wave turbulence at low $R_{e}$ (Tsai et al., 2014; Zhdanov et al., 2015). The present work examines and reports fully developed turbulence through multi-scale vortex dynamics. Our calculations suggest typical values of $R_{e}$ between $8$ and $50$. The Section V appendix includes the calculation of the $R_{e}$. These results suggest dusty plasmas’ similarity to viscoelastic fluids where elastic turbulence is observed at small $R_{e}$ values. This anticipation contradicts presented scalings from energy spectra and structure functions that suggest close alignment with Kolmogorov’s prediction. This remains an open question, and we will try to examine this in future studies.
Supplementary Material
The following information is provided as part of the supplementary material: (I) We provide an agreeable picture of full three-dimensional vortex flow using three different x-z planes at locations $y=-1,0,1$ mm. The energy spectra from all three planes are in the Fig. S1. (II) The movies for campaigns I-VI have been provided in ‘.m4v’ format. (III) The statistical parameters associated with the velocity gradients for campaigns I-VI.
Acknowledgements